If 𝑓 is a continuous function on [0,1], show that lim π‘›β†’βˆž[0 to 1] ∫ 𝑛𝑓(π‘₯)/(1+n^2.π‘₯^2) 𝑑π‘₯ = πœ‹/2. 𝑓(0)


1
Expert's answer
2021-11-03T17:01:49-0400

Let


"I=\displaystyle{\int^{1}_{0}}\frac{nf(x)}{1+n^2x^2}dx=I_1+I_2"


where


"I_1=\displaystyle{\int^{n^{-1/3}}_{0}}\frac{nf(x)}{1+n^2x^2}dx"


and


"I_2=\displaystyle{\int^{1}_{n^{-1/3}}}\frac{nf(x)}{1+n^2x^2}dx"


"|f(x)|\u2264M" Β becauseΒ fΒ is continious. We have:


"I_2\le\displaystyle{\int^{1}_{n^{-1/3}}}|\frac{nf(x)}{1+n^2x^2}|dx\le \displaystyle{\int^{1}_{n^{-1/3}}}\frac{nM}{1+n^2x^2}dx \le \displaystyle{\int^{1}_{n^{-1/3}}}\frac{nM}{1+n^2(n^{-1/3})^2}dx="


"=\frac{nM}{1+n^2(n^{-1/3})^2}(1-n^{-1/3})=o(1)"


Put


"I_3=\displaystyle{\int^{n^{-1/3}}_{0}}\frac{nf(0)}{1+n^2x^2}dx"


We have:


"|I_1-I_3|=|\displaystyle{\int^{n^{-1/3}}_{0}}\frac{n(f(x)-f(0))}{1+n^2x^2}dx|\le"


"\le \displaystyle{\sup_{n\in [0,n^{-1/3}]}}|f(x)-f(0)|\cdot \displaystyle{\int^{n^{-1/3}}_{0}}\frac{ndx}{1+n^2x^2}dx=o(1)\cdot \displaystyle{\int^{n^{-1/3}}_{0}}\frac{d(nx)}{1+n^2x^2}dx="


"=o(1)\cdot arctan(n\cdot n^{-1/3})=o(1)\cdot O(1)"


Hence

"I=I_1+I_2=I_3+(I_1-I_3)+I_2=I_3+o(1)+o(1)"

where


"I_3=f(0)\displaystyle{\int^{n^{-1/3}}_{0}}\frac{d(nx)}{1+n^2x^2}dx=f(0)\cdot arctan(n\cdot n^{-1/3})=f(0)\frac{\pi}{2}+o(1)"


So,


"I\to f(0)\frac{\pi}{2}" as "n\to \infin"